| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  64  
 | 
28 Mar 2024, 14:54 | 
Chan Kim | SPB | Shift summary | X-ray delivery
 
 
 
	12.72 keV, 10 Hz, NOT stable 
delivery, pointing drastically changed in | 
   |  | 
  63  
 | 
23 Mar 2024, 23:12 | 
Peter Zalden | FXE | Shift summary | User experiment transient grating, Cris 
Svetina. (LTP)
 
 
 
	Used 7.11 keV from | 
   |  | 
  62  
 | 
22 Mar 2024, 23:13 | 
Peter Zalden | FXE | Shift summary | User experiment transient grating, Cris 
Svetina. (LTP)
 
 
 
	Used 7.11 keV from | 
   |  | 
  61  
 | 
22 Mar 2024, 06:35 | 
Raphael de Wijn | SPB | Shift summary | X-ray delivery
 
- 9.3 keV, >3 mJ/pulse, beam down 
around 30min | 
   |  | 
  60  
 | 
21 Mar 2024, 23:20 | 
Peter Zalden | FXE | Shift summary | User experiment transient grating, Cris 
Svetina. (LTP)
 
 
 
	Used 7.12 keV from ACC | 
   |  | 
  59  
 | 
21 Mar 2024, 06:33 | 
Raphael de Wijn | SPB | Shift summary | X-ray delivery
 
- 9.3 keV, >3 mJ/pulse, stable 
delivery  | 
   |  | 
  58  
 | 
20 Mar 2024, 23:15 | 
Peter Zalden | FXE | Shift summary | Preparation for user experiment transient 
grating, Cris Svetina
 
 
 
	Used 7.12 keV from ACC | 
   |  | 
  57  
 | 
20 Mar 2024, 23:03 | 
Giuseppe Mercurio | SCS | Shift summary | 
 
Late 
shift summary
 
 
 
	coarse timing
 
	installed 
laser shutter on the laser breadboard and | 
   |  | 
  56  
 | 
20 Mar 2024, 12:41 | 
Frederik Wolff-Fabris | PRC | Issue | BKR called PRC around 06:10AM to inform 
on vacuum issues at SA1 beamline leading 
to a lost beam permission for the North Branch. | 
   |  | 
  55  
 | 
16 Mar 2024, 12:12 | 
Harald Sinn | PRC | Issue | 11:15 SASE1/SPB called that they observed 
and issue with the SASE1 XGM/XTD2: about 
every 35 seconds the power reading jumps | 
   |  | 
  54  
 | 
15 Mar 2024, 23:43 | 
Chan Kim | SPB | Shift summary | **X-ray delivery**
 
 
- 6 keV, 3.2 mJ/pulse, stable delivery
 
 
**Optical laser delivery** | 
   |  | 
  53  
 | 
15 Mar 2024, 18:02 | 
Tommaso Mazza | SQS | Stability Issue | it is not clear why, but the beam position 
on BIU2 has changed over time, even more 
since we checked last. it is off in both | 
     |  | 
  52  
 | 
15 Mar 2024, 07:08 | 
Romain Letrun | SPB | Issue | There are now two watchdogs for the XTD2 
attenuator running at the same time
 
 
 
	SA1_XTD2_WATCHDOG/MDL/ATT1_MONITOR020
 
	SA1_XTD2_OPTICS_PROTECTION/MDL/ATT1
 
 
 
that | 
    |  | 
  51  
 | 
15 Mar 2024, 00:16 | 
Chan Kim | SPB | Shift summary | X-ray delivery
 
 
6 keV, 2.4 mJ/pulse, stable delivery
 
 
Optical laser delivery | 
   |  | 
  50  
 | 
15 Mar 2024, 00:07 | 
Harald Sinn | PRC | Issue | SASE1 bunch number was reduced to 10 pulses 
for no obvious reason. 
 
 
Traced it back to a malfunctioning | 
     |  | 
  49  
 | 
13 Mar 2024, 23:41 | 
Chan Kim | SPB | Shift summary | X-ray delivery
 
 
 
	6 keV, 2.6 mJ/pulse, stable 
delivery
 
 
 
Optical | 
   |  | 
  48  
 | 
13 Mar 2024, 10:09 | 
Romain Letrun | SPB | Status | We checked the XTD2 attenuators following 
the intervention from XRO (compar ewith elog:37, before the intervention). | 
 7x   |  | 
  47  
 | 
05 Mar 2024, 07:41 | 
Marcin Sikorski | SPB | Shift summary | 
 
	10 keV beam
 
	Detector: 
JUNGFRAU in the burst mode (16 memory cells)
 
	Achievements:
 
	
 
		multiple 
sfx data sets (lysozyme) with 4 and 8 pulses | 
   |  | 
  46  
 | 
04 Mar 2024, 09:42 | 
Romain Letrun | SPB | Issue | The number of bunches during the last night 
shift kept being changed down to 24 by the 
device SA1_XTD2_WATCHDOG/MDL/ATT1_MONITOR020 | 
    |  | 
  45  
 | 
04 Mar 2024, 06:52 | 
Raphael de Wijn | SPB | Shift summary | 
 
X-ray 
delivery
 
 
 
	10 keV, up to 128 pulses | 
   |   
 |